OpenStudy (anonymous):

for all integers a,b, and c, if a is congruent to b (mod n), and b is congruent to c (mod n), then a is congruent to c (mod n). Please help prove

OpenStudy (freckles):

This is actually a very common theorem . Try using : a-b=nk b-c=ni where i and k are integers To prove: a-c=nj where j is an integer

OpenStudy (freckles):

have you tried anything yet?

OpenStudy (anonymous):

not yet

OpenStudy (freckles):

if you really want a hint I can tell you but it would be end of the proof basically

OpenStudy (anonymous):

i'd appreciate that hint

OpenStudy (freckles):

try adding the two supposed equations together

OpenStudy (anonymous):

so you would have a-c = n(k + i)

OpenStudy (freckles):

and k+i is an integer since integers are closed under addition

OpenStudy (freckles):

so you do have the a-c=nj which means a is congruent to c mod n

OpenStudy (anonymous):

and that proves the theorem?

OpenStudy (freckles):

yep

OpenStudy (anonymous):

so j = (k+i)?

OpenStudy (freckles):

yes j can be represented as a sum of integers since j itself is an integer

OpenStudy (anonymous):

that makes sense

OpenStudy (zarkon):

I would do it without all the extra letters given n|(a-b) and n|(b-c) then n|(a-b)+(b-c)=a-c

OpenStudy (freckles):

that is only because you are incredible and are data like-minded

OpenStudy (anonymous):

Thanks Zarkon, that was put in a lot simpler terms. Your way helped as well freckles. thanks again